### v1

@@ Abstract
Context. Together with gravity, turbulence, and stellar feedback, magnetic fields (B-fields) are thought to play a critical role in the evolution of molecular clouds and star formation processes. Polarization of thermal dust emission is a popular tracer of B-fields in star-forming regions.
Aims. We aim to map the morphology and measure the strength of B-fields of the nearby molecular cloud, rho Ophiuchus-A ( Oph-A), and then to understand the role of B-fields in regulating star formation and shaping the cloud.
Methods. We have analyzed the far-infrared (FIR) polarization of thermal dust emission observed by SOFIA/HAWC at 89 and 154 m toward the densest part of Oph-A, which is irradiated by the nearby B3/4 star, Oph-S1. These FIR polarimetric maps cover an area of (corresponding to pc2) with an angular resolution of 7.8 and 13.6, respectively.
Results. The Oph-A cloud exhibits well-ordered B-fields with magnetic orientations mainly perpendicular to the ridge of the cloud toward the densest region. We obtain a map of B-field strengths in the range of 0.2-2.5 mG, using the Davis–Chandrasekhar-Fermi method. The B-fields are strongest at the densest part of the cloud, which is associated with the starless core SM1, and decreases toward the outskirts of the cloud. By calculating the map of the mass-to-flux ratio, Alfvén Mach number, and plasma parameter in Oph-A, we find that the cloud is predominantly magnetically sub-critical, sub-Alfvénic, which implies that the cloud is supported by strong B-fields that dominate over gravity, turbulence, and thermal gas energy. Measured B-field strengths at two densest subsregions using other methods that account for the compressible mode are relatively lower than that measured with the DCF method but do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation. A virial analysis suggests that the cloud is gravitationally unbound, which is consistent with the previous detection of numerous starless cores in the cloud. Comparing the magnetic pressure with the radiation pressure from the Oph-S1 star, we find that B-fields are sufficiently strong to support the cloud against radiative feedback and to regulate the shape of the cloud.
@@
B-fields are thought to play a critical role in the star formation process by supporting the material in the cloud against self-gravitational collapse (see, e.g., Mouschovias, 2001; Crutcher, 2012; Pattle & Fissel, 2019; Pattle et al., 2023). The collapse can occur when the dense core is magnetically super-critical, with the ratio of mass to magnetic flux (known as mass-to-flux ratio) greater than one, i.e., when B-fields are not strong enough to prevent the gravitational collapse. The collapse of super-critical cores will lead to the formation of protostars. Conversely, when the core is sub-critical, strong B-fields can act against gravitational collapse. However, several mechanisms might enhance the mass-to-flux ratio and enable the core to collapse, including ambipolar diffusion (see e.g., Mestel & Spitzer, 1956; Mouschovias et al., 2006; Liu et al., 2022) and magnetic reconnection diffusion (Lazarian & Vishniac, 1999; Santos-Lima et al., 2012).
@@
Feedback from massive stars (O- or B-type) can influence the properties of the parental molecular cloud and the star formation processes (e.g., Krumholz et al., 2014; Pabst et al., 2019, 2020). Three-dimensional (3D) magnetohydrodynamic (MHD) simulations have suggested that B-fields also influence the structure of clouds along with radiation pressure feedback from massive stars (e.g., Henney et al., 2009; Mackey & Lim, 2011a). It has been shown that when the B-field orientations are parallel to the ionizing radiation field, the cloud will tend to have a flattened shape, while the cloud will have a column-like or pillar structure if the B-fields are perpendicular to the radiation field (Mackey & Lim, 2011b). These findings from simulations have been in agreement with observations and could explain various globular and elongated substructures in the IC 1396 region (Soam et al., 2018), the famous pillar structure in the M16 region (Pattle et al., 2018), and the Horsehead Nebula (Hwang et al., 2023).
@@
Interstellar dust grains are known to have aspherical shapes and can be systematically aligned with a fixed direction in space, e.g., interstellar B-fields (Hall, 1949; Hiltner, 1949). The alignment of interstellar dust grains with B-fields involves the internal alignment of the grain axis of the maximum inertia moment (also the shortest axis) with its angular momentum (i.e., internal alignment) and the alignment of the angular momentum with B-fields (external alignment). Internal alignment is caused by internal relaxation processes, including Barnett relaxation and inelastic relaxation (Purcell, 1979). The external alignment is caused by the Radiative Torque (RAT) alignment mechanism (Draine & Weingartner, 1997; Lazarian & Hoang, 2007) and by the more comprehensive magnetically-enhanced RAT mechanism (MRAT, Hoang & Lazarian, 2016). The internal and external alignment processes establish the alignment of the grain with its minor axis parallel to the B-fields (see, e.g., Hoang et al., 2022). Dust grains absorb or scatter starlight in the ultraviolet (UV), optical, or near-infrared (NIR) range. However, aligned dust grains tend to absorb/scatter the electric field component of the starlight along their major axis more efficiently. This results in the polarization of background starlight with the polarization vector parallel to the grain’s minor axis. Thus, the polarization of the starlight seen in the UV to optical or NIR wavelength has the polarization orientation parallel to the B-field orientation. Aligned dust grains also re-emit the energy absorbed from starlight in the longer wavelength from far-infrared (FIR) to sub-millimeter (sub-mm), so-called thermal dust emission. As the electric field component of starlight is absorbed more along the grain’s major axis, the dust grain will re-emit more energy with the electric field component along this major axis than along the minor axis. This causes the thermal emission of aligned dust grains to be polarized with the polarization orientation along the grain’s major axis. Because dust grains have their minor axes aligned with the B-field orientation, the polarization orientation of thermal emission is perpendicular to the B-field orientation. Thus, if the position angle of the polarized thermal dust emission is known, one can rotate the polarization angle by 90° to infer the position angle of the B-fields in the medium. Linear polarized thermal dust emission has been routinely used to trace B-fields in various environments of molecular clouds in the Milky Way (see, e.g., Ward-Thompson et al., 2017; Chuss et al., 2019; Pattle et al., 2022) and extra-galactic regions (see, e.g., Lopez-Rodriguez et al., 2021, 2022; Borlaff et al., 2023). The Davis-Chandrasekhar-Fermi method (DCF; Davis, 1951; Chandrasekhar & Fermi, 1953) has been widely used to measure the strengths of the B-field projected in the plane-of-the-sky, , based on the assumption that the kinetic energy of turbulent gas balances to that of turbulent B-fields.
@@
Adopting the mean value of B-field strengths along the line-of-sight () measured using the Zeeman technique over different molecular clouds, Crutcher et al. (2010) derived a relationship between the B-field strengths and gas volume densities (). For low densities of 300 cm-3, B-fields were found not to considerably change with densities. However, for high densities above 300 cm-3, the B-field strengths increase with densities as , where the power-law index is 0.65. Pattle et al. (2023) compiled a large number of measurements of the B-field strength projected in the plane-of-the-sky () estimated using the DCF method and showed that these generally follow Crutcher et al. (2010)’s results until cm-3. Beyond this value, the distribution is more sparse (see Figure 2a in Pattle et al., 2023). Theoretical studies have suggested that the power-law index varies between 2/3 for weak B-fields (Mestel, 1966) and 0.5 for strong B-fields (Mouschovias & Ciolek, 1999). Therefore, the maps of the morphology and strength of the B-fields within a specific cloud are essential to understanding the role of the local B-fields on star formation within the cloud. Previous authors attempted to derive the map of B-fields toward some well-known star-forming regions, for instance, OMC-1 (Guerra et al. 2021, Hwang et al. 2021), 30 Doradus (30 Dor, Tram et al. 2023), and the G11.11–0.12 filament (Ngoc et al. 2023). These studies revealed significant variations in the B-field morphology and strength across the clouds and demonstrated that B-fields strongly influence the cloud’s evolution and, thus, the star formation activities.
@@
In this work, we focus on a specific system of the Ophiuchus-A (hereafter Oph-A) molecular cloud to study the effects of stellar feedback and B-fields on the evolution of the cloud and star formation. The Oph-A molecular cloud is located in the northern part of the L1688 active low-mass star-forming region in the Ophiuchus dark cloud complex (Loren et al., 1990; Wilking et al., 2008; Esplin & Luhman, 2020). The Oph-A cloud was initially identified as a dense gas clump based on DCO+ observations performed by Loren et al. (1990) among other clumps named from Oph-B to Oph-F. This region has high extinctions (20 mag) corresponding to the high gas column densities compared to other regions, which are typically more quiescent (see Figure 1). Being one of the closest low-mass star-forming regions in our Galaxy, located at a distance of 137 pc (Ortiz-León et al., 2017), the gas and dust content in the Oph–A region has been studied extensively at almost every wavelength by both ground-based and space instruments for many years (see e.g., Wilking et al., 2008, for a review). The Oph-A region hosts numerous prestellar and protostellar cores identified by the Ground Belt Survey (Ladjelate et al., 2020) and protostars in the early stages (i.e., Class 0/I Young Stellar Objects (YSOs), Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010). The Oph-A cloud is mostly influenced by a nearby embedded early high-mass B3/4 star (Brown & Zuckerman, 1975; Mookerjea et al., 2018), Oph-S1, which has a mass of 8 (Hamaguchi et al., 2003). The Oph-S1 star is surrounded by a compact H ii region (Andre et al., 1988) and is believed to trigger the star formation activity in Oph–A, especially along the ridge of the dense cloud with indications of protostellar clumps (see Figure 2a in Motte et al., 1998). The star formation activity in the Oph-A is also believed to be affected by the Sco OB2 association located at the western part of the cloud at a distance of 145 pc (de Zeeuw et al., 1999).
@@
B-fields in Oph-A were investigated using data observed in various wavelengths: the NIR (JHK bands) using SIRPOL on the Infrared Survey Facility (IRSF) 1.4–m telescope (Kwon et al., 2015), FIR by the on the 2.7–m Stratospheric Observatory For Infrared Astronomy (SOFIA)/HAWC (Santos et al., 2019), sub-mm by the 15–m James Clerk Maxwell Telescope (JCMT)/POL–2 (Kwon et al., 2018). Santos et al. (2019) performed the analysis on polarimetric data from SOFIA/HAWC observations at 89 and 154 m to study the change in the slope of the polarization spectra with respect to the column densities and dust temperatures in Oph-A. Later, Tram et al. (2021) used a similar SOFIA/HAWC dataset to investigate the relation between the polarization fraction of thermal dust emission versus dust temperatures in the cloud and to understand the mechanism of dust grain alignment and disruption induced by RATs. Kwon et al. (2018) estimated the values in Oph-A using the JCMT/POL–2 data at 850 m toward a few densest parts of the cloud, yielding values in the range of 0.5–5 mG. The highest strength was measured toward the densest part of the cloud associated with the starless core SM1 (see Figure 1). These strengths in Oph-A are comparable to values found in the OMC region (Hwang et al., 2021). Kwon et al. (2018) also found that the B-field direction is generally perpendicular to the main structure of the cloud, which is generally seen in the dense cores or high-density structures of molecular clouds. The Oph-A region provides an ideal isolated system for studying the impact of the B-fields in the star-forming region and investigating the interplay between B-fields and the stellar feedback into the molecular cloud.
In this paper, we aim to:
1.
determine the B-field morphology and strengths in the entire Oph-A region using the SOFIA/HAWC dust polarization data;
2.
investigate the role of B-fields relative to gravity and turbulence in the star formation activity in the cloud;
3.
quantify the combined effect of B-fields and radiation pressure from the nearby Oph-S1 star on the shape of the Oph-A cloud.
@@
The paper is organized as follows. We describe a set of archival observational data in Section 2. We present the results of mapping the B-field morphology and strength using various techniques in Section 3 and discuss the roles of B-fields relative to gravity, turbulence, and radiation pressure feedback from the Oph-S1 star in Section 4. Finally, we summarize our results in Section 5.
@@ 2.1 Dust polarization data from SOFIA/HAWC
We use the FIR polarimetric observations toward Oph-A from the High-resolution Airborne Wideband Camera-plus (HAWC; Harper et al., 2018) onboard SOFIA. The observations were done on 2017 May 17, under the proposal ID (PI: Darren Dowell), using the nod-match-chop mode, with 154° chop angle, 480″ chop throw, and 4-points dithers. Here, we use the SOFIA/HAWC Level–4 data for two bands C and D, centered at 89 and 154 m, respectively, which is fully calibrated and can be directly downloaded from the SOFIA archive111https://irsa.ipac.caltech.edu/Missions/sofia.html. The final polarimetric maps cover an area of toward the central region in Oph-A with the angular resolutions of 7.8 and 13.6 full-width-half-maximum (FWHM) at 89 and 154 m, respectively. The pixel size of HAWC maps is 2.0″ at 89 m and 3.4″ at 154 m. The observations and data reduction are described in detail in Santos et al. (2019).
From the maps of the Stokes-, Stokes-, Stokes- and their uncertainties , , , the maps of polarization fraction () and its corresponding uncertainty () are calculated. The de-biased polarization fraction, , is calculated as: (Vaillancourt, 2006). The maps of polarization angle () and its corresponding uncertainty () in degrees can be derived from Stokes- and Stokes-. We refer the reader to Gordon et al. (2018) for a full description of estimates of the polarization fraction, polarization angle, and their corresponding uncertainties. To enhance the quality in measuring the maps of B-field orientation and strength in Oph-A, we apply some critical conditions in choosing vectors (pixels) in SOFIA/HAWC maps. We mask out from SOFIA/HAWC maps the pixels for which ¡54 for band C and ¡214 for band D. In both bands, we further remove the pixels for which ¡3 and ¿50% from HAWC maps (see Appendix A).

### v2

@@ Abstract
Context. Together with gravity, turbulence, and stellar feedback, magnetic fields (B-fields) are thought to play a critical role in the evolution of molecular clouds and star formation processes. The polarization of thermal dust emission is a popular tracer of B-fields in star-forming regions.
Aims. We aim to map the morphology and measure the strength of B-fields of the nearby molecular cloud, rho Ophiuchus-A ( Oph-A), to understand the role of B-fields in regulating star formation and in shaping the cloud.
Methods. We analyzed the far-infrared (FIR) polarization of thermal dust emission observed by SOFIA/HAWC at 89 and 154 m toward the densest part of Oph-A, which is irradiated by the nearby B3/4 star, Oph-S1. These FIR polarimetric maps cover an area of (corresponding to pc2) with an angular resolution of 7.8 and 13.6, respectively.
Results. The Oph-A cloud exhibits well-ordered B-fields with magnetic orientations that are mainly perpendicular to the ridge of the cloud toward the densest region. We obtained a map of B-field strengths in the range of 0.2-2.5 mG, using the Davis-Chandrasekhar-Fermi (DCF) method. The B-fields are strongest at the densest part of the cloud, which is associated with the starless core SM1, and then decrease toward the outskirts of the cloud. By calculating the map of the mass-to-flux ratio, Alfvén Mach number, and plasma parameter in Oph-A, we find that the cloud is predominantly magnetically sub-critical, sub-Alfvénic, which implies that the cloud is supported by strong B-fields that dominate over gravity, turbulence, and thermal gas energy. The measured B-field strengths at the two densest subsregions using other methods that account for the compressible mode are relatively lower than that measured with the DCF method. However, these results do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation. Our virial analysis suggests that the cloud is gravitationally unbound, which is consistent with the previous detection of numerous starless cores in the cloud. By comparing the magnetic pressure with the radiation pressure from the Oph-S1 star, we find that B-fields are sufficiently strong to support the cloud against radiative feedback and to regulate the shape of the cloud.
@@
B-fields are thought to play a critical role in the star formation process by supporting the material in the cloud against self-gravitational collapse (see, e.g., Mouschovias, 2001; Crutcher, 2012; Pattle & Fissel, 2019; Pattle et al., 2023). The collapse can occur when the dense core is magnetically super-critical, with the ratio of mass to magnetic flux (known as mass-to-flux ratio) greater than 1; that is, when B-fields are not strong enough to prevent the gravitational collapse. The collapse of super-critical cores will lead to the formation of protostars. Conversely, when the core is sub-critical, strong B-fields can act against gravitational collapse. However, several mechanisms might enhance the mass-to-flux ratio and cause the core to collapse, including ambipolar diffusion (see e.g., Mestel & Spitzer, 1956; Mouschovias et al., 2006; Liu et al., 2022) and magnetic reconnection diffusion (Lazarian & Vishniac, 1999; Santos-Lima et al., 2012).
@@
Feedback from massive stars (O- or B-type) can influence the properties of the parental molecular cloud and the star formation processes (e.g., Krumholz et al., 2014; Pabst et al., 2019, 2020). Three-dimensional magnetohydrodynamic (3D MHD) simulations have suggested that B-fields also influence the structure of clouds along with radiation pressure feedback from massive stars (e.g., Henney et al., 2009; Mackey & Lim, 2011a). It has been shown that when the B-field orientations are parallel to the ionizing radiation field, the cloud will tend to have a flattened shape, while the cloud will have a column-like or pillar structure if the B-fields are perpendicular to the radiation field (Mackey & Lim, 2011b). These findings from simulations have been in agreement with observations and could explain various globular and elongated substructures in the IC 1396 region (Soam et al., 2018), the famous pillar structure in the M16 region (Pattle et al., 2018), and the Horsehead Nebula (Hwang et al., 2023).
@@
Interstellar dust grains are known to have aspherical shapes and can be systematically aligned with a fixed direction in space, for instance, interstellar B-fields (Hall, 1949; Hiltner, 1949). The alignment of interstellar dust grains with B-fields involves the internal alignment of the grain axis of the maximum inertia moment (also the shortest axis) with its angular momentum (i.e., internal alignment) and the alignment of the angular momentum with B-fields (external alignment). Internal alignment is caused by internal relaxation processes, including Barnett relaxation and inelastic relaxation (Purcell, 1979). The external alignment is caused by the radiative torque (RAT) alignment mechanism (Draine & Weingartner, 1997; Lazarian & Hoang, 2007) and by the more comprehensive magnetically enhanced RAT mechanism (MRAT, Hoang & Lazarian, 2016). The internal and external alignment processes establish the alignment of the grain with its minor axis parallel to the B-fields (see, e.g., Hoang et al., 2022). Dust grains absorb or scatter starlight in the ultraviolet (UV), optical, or near-infrared (NIR) range. However, aligned dust grains tend to absorb and scatter the electric field component of the starlight along their major axis more efficiently. This results in the polarization of background starlight with the polarization vector parallel to the grain’s minor axis. Thus, the polarization of the starlight seen in the UV to optical or NIR wavelength has the polarization orientation parallel to the B-field orientation. Aligned dust grains also re-emit the energy absorbed from starlight in the longer wavelength from far-infrared (FIR) to sub-millimeter (sub-mm), the so-called thermal dust emission. As the electric field component of starlight is absorbed more along the grain’s major axis, the dust grain will re-emit more energy with the electric field component along this major axis than along the minor axis. This causes the thermal emission of aligned dust grains to be polarized with the polarization orientation along the grain’s major axis. Because dust grains have their minor axes aligned with the B-field orientation, the polarization orientation of thermal emission is perpendicular to the B-field orientation. Thus, if the position angle of the polarized thermal dust emission is known, one can rotate the polarization angle by 90° to infer the position angle of the B-fields in the medium. Linear polarized thermal dust emission has been routinely used to trace B-fields in various environments of molecular clouds in the Milky Way (see, e.g., Ward-Thompson et al., 2017; Chuss et al., 2019; Pattle et al., 2022) and extragalactic regions (see, e.g., Lopez-Rodriguez et al., 2021, 2022; Borlaff et al., 2023). The Davis-Chandrasekhar-Fermi method (DCF; Davis, 1951; Chandrasekhar & Fermi, 1953) has been widely used to measure the strengths of the B-field projected on the plane of the sky, , based on the assumption that the kinetic energy of turbulent gas balances to that of turbulent B-fields.
@@
Adopting the mean value of B-field strengths along the line of sight () measured using the Zeeman technique over different molecular clouds, Crutcher et al. (2010) derived a relationship between the B-field strengths and gas volume densities (). For low densities of 300 cm-3, B-fields were found not to considerably change with densities. However, for high densities above 300 cm-3, the B-field strengths increase with densities as , where the power-law index is 0.65. Pattle et al. (2023) compiled a large number of measurements of the B-field strength projected on the plane of the sky () estimated using the DCF method. These authors showed that these generally follow the results from Crutcher et al. (2010) until cm-3. Beyond this value, the distribution is more sparse (see Fig. 2a in Pattle et al., 2023). Theoretical studies have suggested that the power-law index varies between 2/3 for weak B-fields (Mestel, 1966) and 0.5 for strong B-fields (Mouschovias & Ciolek, 1999). Therefore, the maps of the morphology and strength of the B-fields within a specific cloud are essential to understanding the role of the local B-fields on star formation within the cloud. Previous authors have attempted to derive the map of B-fields toward some well-known star-forming regions, for instance, OMC-1 (Guerra et al., 2021; Hwang et al., 2021), 30 Doradus (30 Dor, Tram et al., 2023), and the G11.11-0.12 filament (Ngoc et al., 2023). These studies revealed significant variations in the B-field morphology and strength across the clouds, demonstrating that B-fields strongly influence the cloud’s evolution and, thus, star formation activities.
@@
In this work, we focus on a specific system of the Ophiuchus-A (hereafter Oph-A) molecular cloud to study the effects of stellar feedback and B-fields on the evolution of the cloud and star formation. The Oph-A molecular cloud is located in the northern part of the L1688 active low-mass star-forming region in the Ophiuchus dark cloud complex (Loren et al., 1990; Wilking et al., 2008; Esplin & Luhman, 2020). The Oph-A cloud was initially identified as a dense gas clump based on DCO+ observations performed by Loren et al. (1990), along with other clumps named Oph-B to Oph-F. This region has high extinctions (20 mag) corresponding to the high gas column densities compared to other regions, which are typically more quiescent (see Fig. 1). Being one of the closest low-mass star-forming regions in our Galaxy, located at a distance of 137 pc (Ortiz-León et al., 2017), the gas and dust content in the Oph–A region has been studied extensively at almost every wavelength by both ground-based and space instruments for many years (see e.g., Wilking et al., 2008, for a review). The Oph-A region hosts numerous prestellar and protostellar cores identified by the Ground Belt Survey (Ladjelate et al., 2020) and protostars in the early stages (i.e., Class 0/I Young Stellar Objects (YSOs), Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010). The Oph-A cloud is mostly influenced by a nearby embedded early high-mass B3/4 star (Brown & Zuckerman, 1975; Mookerjea et al., 2018), Oph-S1, which has a mass of 8 (Hamaguchi et al., 2003). The Oph-S1 star is surrounded by a compact H ii region (Andre et al., 1988) and is believed to trigger the star formation activity in Oph–A, especially along the ridge of the dense cloud with indications of protostellar clumps (see Fig. 2a in Motte et al., 1998). The star formation activity in the Oph-A is also believed to be affected by the Sco OB2 association located in the western part of the cloud at a distance of 145 pc (de Zeeuw et al., 1999).
@@
We investigated B-fields in Oph-A using data observed in various wavelengths: NIR (JHK bands) using imaging polarimeter SIRPOL (Kandori et al., 2006) mounted on the 1.4-m Infrared Survey Facility (IRSF) telescope (Kwon et al., 2015), FIR by the High-resolution Airborne Wideband Camera-plus (HAWC; Harper et al., 2018) on board the 2.7-m Stratospheric Observatory For Infrared Astronomy (SOFIA) (Santos et al., 2019), and sub-mm using the linear polarimeter POL-2 placed in front of the Submillimetre Common-User Bolometer Array 2 (SCUBA-2; Holland et al., 2013) camera mounted on the 15-m James Clerk Maxwell Telescope (JCMT) (Kwon et al., 2018). Santos et al. (2019) performed the analysis on polarimetric data from SOFIA/HAWC observations at 89 and 154 m to study the change in the slope of the polarization spectra with respect to the column densities and dust temperatures in Oph-A. Later, Tram et al. (2021) used a similar SOFIA/HAWC dataset to investigate the relation between the polarization fraction of thermal dust emission versus dust temperatures in the cloud and to understand the mechanism of dust grain alignment and disruption induced by RATs. Kwon et al. (2018) estimated the values in Oph-A, using the JCMT/POL-2 data at 850 m, toward a few of the densest parts of the cloud, yielding values in the range of 0.5–5 mG. The highest strength was measured toward the densest part of the cloud associated with the starless core SM1 (see Fig. 1). These strengths in Oph-A are comparable to values found in the OMC region (Hwang et al., 2021). Kwon et al. (2018) also found that the B-field direction is generally perpendicular to the main structure of the cloud, which is generally seen in the dense cores or high-density structures of molecular clouds. The Oph-A region provides an ideal isolated system for studying the impact of the B-fields in the star-forming region and investigating the interplay between B-fields and the stellar feedback into the molecular cloud. In this paper, we aim to: 1) determine the B-field morphology and strengths in the entire Oph-A region using the SOFIA/HAWC dust polarization data; 2) investigate the role of B-fields relative to gravity and turbulence in the star formation activity in the cloud; and 3) quantify the combined effect of B-fields and radiation pressure from the nearby Oph-S1 star on the shape of the Oph-A cloud.
@@
The paper is organized as follows. We describe a set of archival observational data in Sect. 2. We present the results of mapping the B-field morphology and strength using various techniques in Sect. 3 and discuss the roles of B-fields relative to gravity, turbulence, and radiation pressure feedback from the Oph-S1 star in Sect. 4. Finally, we summarize our results in Sect. 5.
@@ 2.1 Dust polarization data from SOFIA/HAWC
We use the FIR polarimetric observations from the onboard SOFIA/HAWC toward Oph-A. The observations were done on 2017 May 17, under the proposal ID (PI: Darren Dowell), using the nod-match-chop mode, with 154° chop angle, 480″ chop throw, and 4-points dithers. Here, we use the SOFIA/HAWC Level-4 data for two bands C and D, centered at 89 and 154 m, respectively, which is fully calibrated and can be directly downloaded from the SOFIA archive111https://irsa.ipac.caltech.edu/Missions/sofia.html. The final polarimetric maps cover an area of (corresponding to pc2) toward the central region in Oph-A with the angular resolutions of 7.8 and 13.6 full-width-half-maximum (FWHM) at 89 and 154 m, respectively. The pixel size of HAWC maps is 2.0″ at 89 m and 3.4″ at 154 m. The observations and data reduction are described in detail in Santos et al. (2019).
From the maps of the Stokes-, Stokes-, and Stokes- (along with their uncertainties , , ), the maps of polarization fraction () and its corresponding uncertainty () were calculated. The de-biased polarization fraction, , is calculated as: (Vaillancourt, 2006). The maps of polarization angle () and its corresponding uncertainty () in degrees can be derived from Stokes- and Stokes-. We refer to Gordon et al. (2018) for a full description of estimates of the polarization fraction, polarization angle, and their corresponding uncertainties. To enhance the quality in measuring the maps of B-field orientation and strength in Oph-A, we apply some critical conditions in choosing vectors (pixels) in SOFIA/HAWC maps. We mask out from SOFIA/HAWC maps the pixels for which ¡54 for band C and ¡214 for band D. In both bands, we further removed the pixels for which we have ¡3 and ¿50% from the HAWC maps (see Appendix A).